OpenStudy (anonymous):

Mrs. jones puts 3 black marbles, 4 white marbles, and 3 red marbles in a jar. What is the probability that the first two marbles she draws out randomly will be red and white? A 7/20 B 7/19 C 3/25 D 2/15 @sourwing

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

I'm not good with probability, but based off what sourwing did before... 3+4+3=10 10*10=100 3*4=12 reduce fraction 12/100 = ?

OpenStudy (anonymous):

3/25 thanks!